# Harnessing the Power of Neural Operators with Automatically Encoded Conservation Laws

### v2

@@ 3.2 ClawNO architecture and implementation
Figure 2: Illustration of the proposed clawNO architecture. We start from the input function $\mathbf{g}(x)$. After lifting, the high-dimensional latent representation goes through a series of iterative equivariant layers, then gets projected to a function space in the form of an antisymmetric matrix. Lastly, we employ numerical differentiation (layer $\mathcal{D}$ with pre-calculated weights) to obtain the target divergence-free output.
With the above analysis, we now construct an NO architecture with divergence-free output. Concretely, we propose to modify the architecture in equation 2 as:
$\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x}):=\mathcal{D}\circ\mathcal{Q}\circ\mathcal{J}_{L}\circ\cdots\circ\mathcal{J}_{1}\circ\mathcal{P}[\mathbf{f}](\mathbf{x}),\vspace{-0.03in}$ (6)
@@
To formulate the last layer $\mathcal{D}$ there remains one more technical challenge: in NOs the output functions are provided on a set of discrete grid points, $\chi$, and therefore the derivatives are not readily evaluated. With the key observation that in classical numerical differentiation methods the derivatives are evaluated as weighted linear combinations of neighboring or global points, we propose to encode these weights in $\mathcal{D}$. As such, $\mathcal{D}$ will have pre-calculated weights and act as a numerical approximation of $[\text{div}(\mu_{1}),\cdots,\text{div}(\mu_{p})]^{T}$. To demonstrate the idea, in the following we start with uniform grids $\chi$ and ${\Omega}$ being a periodic piece of $\mathbb{R}^{p}$, i.e., the torus $\mathbb{T}^{p}$, then discuss the scenarios with non-periodic domain and non-uniform grids.
@@
$\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}(\mathbf{x}):=\sum_{\xi_{1}=-N^{(1)}/2}^{N^{(1)}/2-1}\cdots\sum_{\xi_{p}=-N^{(p)}/2}^{N^{(p)}/2-1}\hat{\mu}_{jk}(\xi_{1},\cdots,\xi_{p})e^{i2\pi\xi_{1}/L^{(1)}}\cdots e^{i2\pi\xi_{p}/L^{(p)}},\vspace{-0.03in}$
@@
$(\mathcal{D}[\mu])_{j}=\sum_{k}\dfrac{\partial\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}}{\partial x_{k}}=\sum_{k}\mathcal{F}^{-1}\left[i\frac{2\pi\xi_{k}}{L^{(k)}}\mathcal{F}[\mu_{jk}]({\bm{\xi}})\right],\vspace{-0.05in}$ (7)
@@
${\left|\left|\star d\mu-\mathcal{D}[\mu]\right|\right|}_{L^{\infty}}=O(\Delta x^{m-1}),\vspace{-0.05in}$
@@
Proof. Derivation of equation 7 is provided in Appendix D.1. The theorem is an immediate proposition of the Fourier spectral differentiation error estimate (see, e.g., Trefethen (2000, Page 34)).∎
We now consider non-periodic domain and/or non-uniform grids. When the domain is non-periodic but the grids are uniform, we employ the Fourier continuation (FC) technique (Amlani & Bruno, 2016; Maust et al., 2022) which extends the non-periodic model output into a periodic function. In particular, the layer parameter for $\mathcal{D}$ in equation 7 is modified as: $\tilde{\mathcal{D}}:=\mathcal{R}\circ\mathcal{D}\circ\textbf{FC}$, where FC and $\mathcal{R}$ are the FC extension and restriction operators, respectively.
@@
$\min_{\omega^{(k)}_{i,j}}\sum_{\mathbf{x}_{j}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}|\omega^{(k)}_{i,j}|,\;\text{s.t.}\;\frac{\partial\phi}{\partial x_{k}}(\mathbf{x}_{i})=\sum_{\mathbf{x}_{j}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}(\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i}))\omega^{(k)}_{i,j},\quad\forall\phi\in\mathbb{P}_{m}(\mathbb{R}^{p}),\vspace{-0.03in}$ (8)
where $\mathbb{P}_{m}(\mathbb{R}^{p})$ is the space of $m-$th order polynomials ($m\geq 1$). Assume that $\chi$ is quasi-uniform, the size of $\chi\cap B_{\delta}(\mathbf{x})$ is bounded, the domain ${\Omega}$ satisfies a cone condition, and $\delta$ is sufficiently large, the above optimization problem has a solution (Wendland, 2004), and we further have:
@@
$(\mathcal{D}[\mu])_{j}(\mathbf{x}_{i})=\sum_{k}\sum_{\mathbf{x}_{l}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}(\mu_{jk}(\mathbf{x}_{l})-\mu_{jk}(\mathbf{x}_{i}))\omega^{(k)}_{i,l}.\vspace{-0.1in}$ (9)
@@
${\left|\left|\star d\mu-\mathcal{D}[\mu]\right|\right|}_{L^{\infty}}=O(\Delta x^{m+1}),\vspace{-0.05in}$
@@
Proof. The proof follows a similar argument as in Levin (1998, Theorem 5), see Appendix D.2. ∎
The proposed architecture can be readily combined with the FNOs in equation 3 and GNOs in equation 4. Generally, when measurements are provided on regular domain and uniform grids, we consider FNOs together with the differentiation layer $\mathcal{D}$ approximated in equation 7 (or its FC-based extension). When the measurements are on non-uniform grids, GNOs can be employed with the differentiation layer $\mathcal{D}$ approximated via meshfree methods (see equation 9). Besides vanilla FNOs and GNOs, to further encode laws of physics we also consider their group equivariant versions as the base models. Namely, the invariant neural operator (INO) ((Liu et al., 2023)) for non-uniform grids and the group-equivariant FNO (G-FNO) ((Helwig et al., 2023)) for uniform grids. As such, the learnt laws of physics do not depend on the coordinate system used to describe them, together with the encoded conservation laws guaranteed.
@@ 4 Experiments
We showcase the prediction accuracy and expressivity of clawNOs across a wide range of scientific problems, including elasticity, shallow water equations and incompressible Navier-Stokes equations. We compare the performance of clawNOs against a number of relevant machine learning techniques. In particular, we select INO (Liu et al., 2023) and GNO (Li et al., 2020a) as baselines in graph-based settings, whereas we choose FNO (Li et al., 2020c), G-FNO (Helwig et al., 2023), UNet (Ronneberger et al., 2015), LSM (Wu et al., 2023b), KNO (Xiong et al., 2023a; b), and UNO (Rahman et al., 2022) as baselines in datasets with uniform discretizations. We also compare with an additional non-NO baseline, Lagrangian neural networks (LNN) (Cranmer et al., 2020; Müller, 2023), in the first experiment. The relative L2 error is reported as comparison metrics. We perform three replicates of all the experiments with randomly selected seeds, and the mean and standard deviation of the errors are reported. To guarantee a fair comparison, we make sure that the total numbers of trainable parameters in all models are on the same level, and report these numbers for each case. Additional details on the data generation and training strategies are provided in Appendix B.
@@ 4.1 Incompressible Navier-Stokes equation
$\nabla\cdot\mathbf{u}=0\text{ ,}\quad\partial_{t}w+\mathbf{u}\cdot\nabla w=\nu\Delta w+\mathbf{f}\text{ ,}\vspace{-0.05in}$ (10)
@@ 4.2 Radial dam break
$\partial_{t}h+\nabla\cdot\left(h\mathbf{u}\right)=0\text{ ,}\quad\partial_{t}\left(h\mathbf{u}\right)+\nabla\cdot\left(h\mathbf{u}\mathbf{u}^{T}\right)+\frac{1}{2}g\nabla h^{2}+gh\nabla b=0\text{ ,}\vspace{-0.05in}$ (11)
@@ 4.3 Atmospheric modeling
$\frac{\partial\zeta}{\partial t}+\nabla\cdot(\mathbf{u}(\zeta+f))=0\text{ ,}\quad\frac{\partial\mathcal{D}}{\partial t}-\nabla\times(\mathbf{u}(\zeta+f))=-\nabla^{2}(\tfrac{1}{2}|\mathbf{u}|^{2}+g\eta)\text{ .}\vspace{-0.05in}$ (12)
@@
We list in Table 3 our experimental results along with comparisons against selected baselines. For this problem, since the spherical coordinate system in latitudes and longitudes stay the same and thus the equivariance property plays little role, we directly employ clawFNO as our representative clawNO model, as the latent dimension of clawFNO is much higher than that of clawGFNO for similar total numbers of model parameters. In other words, clawFNO possesses more expressivity power. This is rather important in atmospheric modeling at the entire globe scale, as the wave propagation patterns are more localized compared to other examples. In this context, clawFNO outperforms the best non-claw baseline by 47% in accuracy. Additionally, it is worth mentioning that the majority of the errors accumulate in the vicinity of mountains, as evidenced in the rollout demonstration in Figures 11-13 (compare to Fig. 5a). This is expected as the interaction with the orography causes wave reflection and dispersion through changes in the wave speed, which generally limits predictability more than in other regions. Furthermore, the training strategy (Appendix B.5) disregards wave numbers higher than 22 which, however, are more pronounced around rough orography. Our simulations resolve orography up to wave number 63 (Appendix B.3).
4.4 Constitutive modeling of material deformation
In this example we test the efficacy of the proposed clawNO in graph-based settings. Specifically, we consider the elastic deformation in materials, where the (hidden) constitutive law reads:
$-\nabla\cdot\sigma=\mathbf{f}\text{, }\quad\sigma=\lambda(\nabla\cdot\mathbf{u})\mathbf{I}+\mu(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})\text{, }\quad\mathbf{u}|_{\partial\Omega}=\mathbf{u}_{D}\text{, }\vspace{-0.05in}$ (13)
with Dirichlet boundary conditions, and the (known) mass conservation law $\nabla\cdot\mathbf{u}=0$ imposed. Here $\sigma$ is the stress tensor, $\mathbf{u}$ denotes the displacement, $\mathbf{f}$ denotes the body load, $\lambda$ and $\mu$ are the Lame constants. Under this setting, the Dirichlet boundary condition and the body load are treated as inputs. To investigate the efficacy on irregular grids, we consider a circular domain of radius $0.4$.
Table 4: Test errors and the number of trainable parameters for the material deformation problem, where bold numbers highlight the best method.
@@ 4.4 Constitutive modeling of material deformation
clawINO 4,747,521 8.54%$\pm$0.125.88%$\pm$0.043.92%$\pm$0.16%
INO 4,747,521 10.51%$\pm$0.616.66%$\pm$0.254.11%$\pm$0.64%
GNO 18,930,498 10.02%$\pm$0.267.68%$\pm$0.055.82%$\pm$0.46%
We report in Table 4 the numerical results of clawINO, along with the comparison against GNO and INO in various data regimes. Consistent with the findings on regular grids, clawINO outperforms other graph-based models in all data regimes, beating the best non-claw baseline by 14.8%, 11.7%, and 4.6% in the considered three data regimes, respectively. Note that the total number of parameters in GNO is roughly four times the number of parameters in INO-based models, as the parameters in INO-based models are layer-independent, which also results in an advantage in memory cost.
@@ 5 Conclusion
We introduce a series of clawNOs that explicitly bake fundamental conservation laws into the neural operator architecture. In particular, we build divergence-free output functions in light of the concept of differential forms and design an additional layer to recover the divergence-free target functions based on numerical differentiation. We perform extensive experiments covering a wide variety of challenging scientific machine learning problems, and show that it is essential to encode conservation laws for correct physical realizability, especially in small-data regimes. For the future work we will explore more fundamental laws in other application domains. We also point out that the idea of incorporating numerical PDE techniques into NOs can also be extended. For instance, the quadrature weights can provide a more accurate numerical integration in evaluating the cost functional, as compared to the current mean squared error widely employed in NOs.

### v3

@@ 3.2 ClawNO Architecture and Implementation
Figure 2: Proposed clawNO architecture. We start from the input function $\mathbf{g}(x)$. After lifting, the high-dimensional latent representation goes through a series of iterative equivariant layers, then gets projected to a function space in the form of an antisymmetric matrix. Lastly, we employ numerical differentiation (layer $\mathcal{D}$ with pre-calculated weights) to obtain the target divergence-free output.
With the above analysis, we now construct an NO architecture with divergence-free output. Concretely, we propose to modify the architecture in (2) as:
$\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x}):=\mathcal{D}\circ\mathcal{Q}\circ\mathcal{J}_{L}\circ\cdots\circ\mathcal{J}_{1}\circ\mathcal{P}[\mathbf{f}](\mathbf{x})\text{ ,}$ (6)
@@
To formulate the last layer $\mathcal{D}$ there remains one more technical challenge: in NOs the output functions are provided on a set of discrete grid points, $\chi$, and therefore the derivatives are not readily evaluated. With the key observation that in classical numerical differentiation methods the derivatives are evaluated as weighted linear combinations of neighboring or global points, we propose to encode these weights in $\mathcal{D}$. As such, $\mathcal{D}$ will have pre-calculated weights and act as a numerical approximation of $[\text{div}(\mu_{1}),\cdots,\text{div}(\mu_{p})]^{T}$. To demonstrate the idea, in the following we start with uniform grids $\chi$ and ${\Omega}$ being a periodic piece of p, i.e., the torus $\mathbb{T}^{p}$, then discuss the scenarios with non-periodic domain and non-uniform grids.
@@
$\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}(\mathbf{x}):=$
$\sum_{\xi_{1}=-\frac{N^{(1)}}{2}}^{\frac{N^{(1)}}{2}-1}\cdots\sum_{\xi_{p}=-\frac{N^{(p)}}{2}}^{\frac{N^{(p)}}{2}-1}\hat{\mu}_{jk}(\xi_{1},\cdots,\xi_{p})e^{\frac{i2\pi\xi_{1}}{L^{(1)}}}\cdots e^{\frac{i2\pi\xi_{p}}{L^{(p)}}}\text{ ,}$
@@
$(\mathcal{D}[\mu])_{j}=\sum_{k}\dfrac{\partial\mu^{(N^{(1)},\cdots,N^{(p)})}_{jk}}{\partial x_{k}}=\sum_{k}\mathcal{F}^{-1}\left[i\frac{2\pi\xi_{k}}{L^{(k)}}\mathcal{F}[\mu_{jk}]({\boldsymbol{\xi}})\right],$ (7)
@@
${\left|\left|\star d\mu-\mathcal{D}[\mu]\right|\right|}_{L^{\infty}}=O(\Delta x^{m-1})\text{ ,}$
@@
Proof. Derivation of (7) is provided in Appendix E.1. The theorem is an immediate proposition of the Fourier spectral differentiation error estimate (see, e.g., Trefethen (2000, Page 34)).∎
We now consider non-periodic domain and/or non-uniform grids. When the domain is non-periodic but the grids are uniform, we employ the Fourier continuation (FC) technique (Amlani & Bruno, 2016; Maust et al., 2022) which extends the non-periodic model output into a periodic function. In particular, the layer parameter for $\mathcal{D}$ in (7) is modified as: $\tilde{\mathcal{D}}:=\mathcal{R}\circ\mathcal{D}\circ\textbf{FC}$, where FC and $\mathcal{R}$ are the FC extension and restriction operators, respectively.
@@
$\min_{\omega^{(k)}_{i,j}}\sum_{\mathbf{x}_{j}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}|\omega^{(k)}_{i,j}|,\;\text{s.t.}\;\frac{\partial\phi}{\partial x_{k}}(\mathbf{x}_{i})$
$=$ $\sum_{\mathbf{x}_{j}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}(\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i}))\omega^{(k)}_{i,j},\quad\forall\phi\in\mathbb{P}_{m}(^{p})\text{ ,}$ (8)
where $\mathbb{P}_{m}(^{p})$ is the space of $m-$th order polynomials ($m\geq 1$). Assume that $\chi$ is quasi-uniform, the size of $\chi\cap B_{\delta}(\mathbf{x})$ is bounded, the domain ${\Omega}$ satisfies a cone condition, and $\delta$ is sufficiently large, the above optimization problem has a solution (Wendland, 2004), and we further have:
@@
$(\mathcal{D}[\mu])_{j}(\mathbf{x}_{i})=\sum_{k}\sum_{\mathbf{x}_{l}\in\chi\cap B_{\delta}(\mathbf{x}_{i})}(\mu_{jk}(\mathbf{x}_{l})-\mu_{jk}(\mathbf{x}_{i}))\omega^{(k)}_{i,l}.$ (9)
@@
${\left|\left|\star d\mu-\mathcal{D}[\mu]\right|\right|}_{L^{\infty}}=O(\Delta x^{m+1})\text{ ,}$
@@
Proof. The proof follows a similar argument as in Levin (1998, Theorem 5), see Appendix E.2. ∎
The proposed architecture can be readily combined with the FNOs in (3) and GNOs in (4). Generally, when measurements are provided on regular domain and uniform grids, we consider FNOs together with the differentiation layer $\mathcal{D}$ approximated in (7) (or its FC-based extension). When the measurements are on non-uniform grids, GNOs can be employed with the differentiation layer $\mathcal{D}$ approximated via meshfree methods (see (9)). Besides vanilla FNOs and GNOs, to further encode laws of physics we also consider their group-equivariant versions as the base models, namely the invariant neural operator (INO) (Liu et al., 2023) for non-uniform grids and the group-equivariant FNO (G-FNO) (Helwig et al., 2023) for uniform grids. As such, the learnt laws of physics do not depend on the coordinate system used to describe them, together with the encoded conservation laws guaranteed.
Remark: Although clawNOs are designed to make divergence-free predictions, they are also applicable to problems involving a mixture of divergence-free and non-divergence-free fields. Assume the number of divergence-free and non-divergence-free variable dimensions are $n_{1}$ and $n_{2}$, respectively. In this situation, one can set the output dimension of the projection layer to be $n=n_{1}+n_{2}$, and only pass the first $n_{1}$ dimensions of the output to the numerical differentiation layer for prediction of divergence-free output, and the remaining $n_{2}$ dimensions can be directly trained to predict the non-divergence-free variables in the spirit of a vanilla NO.
@@ 4 Experiments
We showcase the prediction accuracy and expressivity of clawNOs across a wide range of scientific problems, including elasticity, shallow water equations and incompressible Navier-Stokes equations. We compare the performance of clawNOs against a number of relevant machine learning techniques. In particular, we select INO (Liu et al., 2023), GNO (Li et al., 2020a), MP-PDE (Brandstetter et al., 2022), and MeshGraphNets (Pfaff et al., 2020) as baselines in graph-based settings, whereas we choose FNO (Li et al., 2020c), G-FNO (Helwig et al., 2023), UNet (Ronneberger et al., 2015), LSM (Wu et al., 2023b), KNO (Xiong et al., 2023a, b), UNO (Rahman et al., 2022), and BOON (Saad et al., 2022) as baselines in datasets with uniform discretizations. We also compare with an additional non-NO baseline, Lagrangian neural networks (LNN) (Cranmer et al., 2020; Müller, 2023), in the first experiment. The relative L2 error is reported as comparison metrics. We perform three replicates of all the experiments with randomly selected seeds, and the mean and standard deviation of the errors are reported. To guarantee a fair comparison, we make sure that the total numbers of trainable parameters in all models are on the same level, and report these numbers for each case. Additional details on the data generation and training strategies are provided in Appendix C.
@@ 4.1 Incompressible Navier-Stokes Equation
$\nabla\cdot\mathbf{u}=0\text{ ,}\quad\partial_{t}w+\mathbf{u}\cdot\nabla w=\nu\Delta w+\mathbf{f}\text{ ,}$ (10)
@@ 4.2 Radial Dam Break
$\partial_{t}h+\nabla\cdot\left(h\mathbf{u}\right)$ $=0\text{ ,}$
$\partial_{t}\left(h\mathbf{u}\right)+\nabla\cdot\left(h\mathbf{u}\mathbf{u}^{T}\right)+\frac{1}{2}g\nabla h^{2}+gh\nabla b$ $=0\text{ ,}$ (11)
@@ 4.3 Atmospheric Modeling
[…]
$\frac{\partial\mathcal{D}}{\partial t}-\nabla\times(\mathbf{u}(\zeta+f))=$ $-\nabla^{2}(\tfrac{1}{2}|\mathbf{u}|^{2}+g\eta)\text{ .}$ (12)
@@
Table 4: Test errors and the number of trainable parameters for the material deformation problem. Bold numbers highlight the best method.
@@
clawINO 4,747,521 14.12%$\pm$0.235.41%$\pm$0.671.00%$\pm$0.02%
INO 4,747,521 16.75%$\pm$1.247.69%$\pm$0.512.40%$\pm$0.53%
GNO 18,930,498 19.73%$\pm$3.6012.59%$\pm$0.248.51%$\pm$5.29%
MP-PDE 18,368,610 23.07%$\pm$20.3912.38%$\pm$0.411.30%$\pm$0.55%
MeshGraphNets 18,694,914 38.57%$\pm$17.847.35%$\pm$0.541.79%$\pm$0.81%
We list in Table 3 our experimental results along with comparisons against selected baselines. For this problem, since the spherical coordinate system in latitudes and longitudes stay the same and thus the equivariance property plays little role, we directly employ clawFNO as our representative clawNO model, as the latent dimension of clawFNO is much higher than that of clawGFNO for similar total numbers of model parameters. In other words, clawFNO possesses more expressivity power. This is rather important in atmospheric modeling at the entire globe scale, as the wave propagation patterns are more localized compared to other examples. In this context, clawFNO outperforms the best non-claw baseline by 47% in accuracy. Additionally, it is worth mentioning that the majority of the errors accumulate in the vicinity of mountains, as evidenced in the rollout demonstration in Figures 13-15 (compare to Fig. 6a). This is expected as the interaction with the orography causes wave reflection and dispersion through changes in the wave speed, which generally limits predictability more than in other regions. Furthermore, the training strategy (Appendix C.5) disregards wave numbers higher than 22 which, however, are more pronounced around rough orography. Our simulations resolve orography up to wave number 63 (Appendix C.3).
4.4 Constitutive Modeling of Material Deformation
In this example we test the efficacy of the proposed clawNO in graph-based settings. Specifically, we consider the incompressible material deformation of the Mooney–Rivlin type, where the (hidden) constitutive law reads:
$-\nabla\cdot\sigma$ $=\mathbf{f}\text{ , where }\sigma=-p\mathbf{I}+2C_{1}\mathbf{B}-2C_{2}\mathbf{B}^{-1}\text{ ,}$ (13)
$\mathbf{u}|_{\partial\Omega}$ $=\mathbf{u}_{D}\text{ ,}$
with Dirichlet boundary conditions, and the (known) mass conservation law $\nabla\cdot\mathbf{u}=0$ imposed. Here $\sigma$ is the stress tensor, $\mathbf{u}$ denotes the displacement, $\mathbf{f}$ denotes the body load, $\lambda$ and $\mu$ are the Lame constants. $\mathbf{B}$ is the left Cauchy–Green deformation tensor, $\mathbf{B}:=\mathbf{F}\mathbf{F}^{T}\approx\mathbf{I}+2\text{sym}(\nabla\mathbf{u})$ for infinitesimal strains. Under this setting, the Dirichlet boundary condition and the body load are treated as inputs. To investigate the efficacy on irregular grids, we consider a circular domain of radius $0.4$. We set the material constants $C_{1}=C_{2}=0.075$, and the body load $\mathbf{f}$ is randomly generated following:
$\mathbf{f}(x,y)=\sum_{k_{1},k_{2}=1}^{2}A_{k_{1},k_{2}}\sin(k_{1}\pi x)\sin(k_{2}\pi y)\text{, }$
$A_{k_{1},k_{2}}=\exp(-0.1k_{1}k_{2})\xi\text{ ,}\quad\xi\sim\mathcal{U}[-2.5,2.5]\text{ .}$
We report in Table 4 the numerical results of clawINO, along with the comparison against GNO, INO, MP-PDE and MeshGraphNets in various data regimes. Consistent with the findings on regular grids, clawINO outperforms other graph-based models in all data regimes, beating the best non-claw baseline by 15.7%, 26.4%, and 23.1% in the considered three data regimes, respectively. Note that the total number of parameters in GNO is roughly four times the number of parameters in INO-based models, as the parameters in INO-based models are layer-independent, which also results in an advantage in memory saving.
@@ 5 Conclusion
We introduce a series of clawNOs that explicitly bake fundamental conservation laws into the neural operator architecture. In particular, we build divergence-free output functions in light of the concept of differential forms and design an additional layer to recover the divergence-free target functions based on numerical differentiation. We perform extensive experiments covering a wide variety of challenging scientific machine learning problems, and show that it is essential to encode conservation laws for correct physical realizability, especially in small-data regimes.
For future work, we plan to explore more fundamental laws in other application domains. In some applications such as the shock capturing problem (Hansen et al., 2023), the solution regularity deteriorates and error estimates in this work do not apply. It would be interesting to modify the differentiation layer in clawNOs based on discontinuity capturing methods such as the Essentially Non-Oscillatory method (Shu, 1999) and the discontinuous Galerkin method (Huerta et al., 2012). We also point out that the idea of incorporating numerical PDE techniques into NOs can also be extended. For instance, the quadrature weights can provide a more accurate numerical integration in evaluating the cost functional, as compared to the current mean squared error widely employed in NOs.
Acknowledgements
Y. Fan, L. Zhang, and Y. Yu would like to acknowledge support by the National Science Foundation under award DMS-1753031 and the AFOSR grant FA9550-22-1-0197. Portions of this research were conducted on Lehigh University’s Research Computing infrastructure partially supported by NSF Award 2019035. The authors would also like to thank the reviewers for the careful reading and valuable suggestions that helped improve the quality of the paper.
Impact Statement
Broader Impact: Conservation laws are fundamentally properties of nature, describing that a particular measurable property of a physical system does not change as the system evolves. In physics-based machine learning applications such as material modeling and climate science, the underlying governing equations are often unknown, while the system should satisfy some fundamental physical laws such as the conservation of mass, momentum and energy. Our work focuses on developing a novel neural network architecture to impose these fundamental physical laws, and therefore contributes mostly to the field of scientific machine learning, especially in the enduring problem of data-driven model discovery from physics-based measurements.
Ethics Statement: The data utilized in our study is solely sourced from publicly available datasets or softwares. No user information was involved. We are committed to upholding the principles of responsible and transparent data usage throughout our work.